# Breaking the SFT Plateau: Multimodal Structured Reinforcement Learning for Chart-to-Code Generation

### v1

@@ Abstract
While reinforcement learning (RL) has proven highly effective for general reasoning in vision-language models, its application to tasks requiring in-depth understanding of information-rich images and the generation of structured outputs remains underexplored. Chart-to-code generation is a prime example of this challenge, as it demands complex reasoning over visual charts to generate structured code. Therefore, supervised fine-tuning (SFT) alone is often insufficient, highlighting a critical need for effective RL strategies that appropriately reward the structured outputs. In this paper, we systematically investigate the performance plateau in SFT through large-scale experiments and propose Multimodal Structured Reinforcement Learning (MSRL) for chart-to-code generation, which substantially breaks through this plateau. To this end, we first construct the largest training corpus to date, containing 3 million chart-code pairs. To mitigate the simplistic patterns of prior synthetic data, we curate real-world tables from arXiv papers as the data source. Despite reaching a state-of-the-art performance, our experiments demonstrate that simply scaling the SFT data eventually hits a performance plateau where further increases yield negligible improvements. To break this plateau, our MSRL method leverages a multi-granularity structured reward system that uses multimodal textual and visual feedback to guide the RL training. At the textual level, rule-based rewards validate the fidelity of fine-grained code details. At the visual level, a model-based reward is derived by first rendering the generated code into an image and then employing an evaluator model to assess its structural similarity to the ground-truth chart. To ensure training stability, we implement this multimodal reward within a two-stage curriculum, where the model initially trains on textual rewards alone before progressing to the combined textual and visual signals. Experimental results demonstrate that our MSRL framework significantly breaks the SFT plateau, improving the high-level metrics by 6.2% and 9.9% on ChartMimic and ReachQA benchmarks respectively. Notably, our method significantly outperforms all methods in the chart domain and achieves competitive performance with advanced closed-source models for the first time.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive reasoning capabilities on complex textual problems, including code generation and mathematical problem-solving OpenAI (2025b); Guo et al. (2025). Pivotal to this success is the implementation of reinforcement learning (RL) to optimize the capacities of LLM to generate step-by-step reasoning Wei et al. (2022) and verify final answers Shao et al. (2024). While multimodal large language models (MLLMs) have achieved impressive performance on established visual reasoning tasks like visual question answering Zhang et al. (2024a); Huang et al. (2025), their core reasoning capabilities remain limited, especially when tasks requiring processing information-dense images like charts Xu et al. (2024); Chen et al. (2025a), and generating structured outputs like code Xia et al. (2024); Zhao et al. (2025).
Chart-to-code generation, a complex task characterized by dense visual inputs and structured code outputs, has recently garnered significant attention Shi et al. (2024); Zhang et al. (2024c). Although previous works have attempted to solve this task by using methods like Supervised Fine-Tuning (SFT) Zhang et al. (2024b); Zhao et al. (2025) or Direct Preference Optimization (DPO) Rafailov et al. (2023); Zhang et al. (2025) on synthetic data, their performance remains limited due to the simplistic patterns of the corpus and ineffective generalization strategies. Therefore, a promising direction is to leverage more realistic data and robust strategies to enhance model performance.
@@
In this work, we first systematically investigate the SFT performance plateau through experiments at various data scales. Based on this analysis, we propose a multimodal structured reinforcement learning (MSRL) strategy designed to optimize the model at multiple granularities. Specifically, to investigate the performance plateau of SFT, we first construct the largest training corpus to date, which covers 24 distinct chart types and comprises 3 million chart-code pairs. To ensure the charts reflect realistic distributions and avoid monotonous trends, we design a two-stage generation pipeline that curates high-quality tables from arXiv papers and subsequently leverages these data to generate diverse and complex plot codes utilizing LLMs. Unlike previous works that apply RL directly after SFT Ni et al. (2025); Chen et al. (2025b), our approach first investigates the performance plateau of SFT models. Establishing this baseline is crucial for isolating the true effectiveness of the MSRL strategy from gains that merely compensate for an unsaturated SFT model. To this end, we perform SFT on models with various data scales and reveal that simply increasing the SFT data volume leads to a performance plateau, which cannot be overcome by data quantity alone. To break through this plateau, we propose the MSRL training strategy, which incorporates multi-granularity reward functions from both textual and visual aspects to optimise the model. For the textual aspect, to handle format diversity in code generation, we first filter our training corpus to enforce a consistent structure, making it suitable for information extraction. Building on this standardized data, we develop a comprehensive reward function that verifies code correctness from five key perspectives. However, a reward function based solely on text is often insufficient, as it struggles to capture the overall structures of charts. To address this limitation, we introduce a visual reward mechanism. The core of this method is a render-and-compare process that involves rendering the generated code into an image and then employing an MLLM to score the visual similarity between the rendered image and the original input chart. By leveraging its multi-granularity reward function, MSRL optimizes both global context and fine-grained details. We evaluate models on various chart-to-code benchmarks Shi et al. (2024); He et al. (2024). The results demonstrate that it establishes a new state-of-the-art, as it outperforms all previous open-source models and rivals the performance of powerful proprietary MLLMs. In summary, the main contributions of this work are as follows:
@@ 2.1 Chart MLLMs
Chart understanding is a crucial research area encompassing both low-level tasks, such as data extraction Liu et al. (2022), and high-level tasks like question answering Masry et al. (2022) and summarization Kantharaj et al. (2022). Recent work has focused on training MLLMs on extensive, chart-specific datasets to enhance their understanding capabilities Xia et al. (2024); Zhang et al. (2024b), leading to superior performance across various chart-related tasks. For example, ChartMoE Xu et al. (2024) proposes a mixture-of-experts (MoE) structure for multi-task pretraining. More recently, leveraging reinforcement learning to enhance reasoning capacities has garnered significant interest. Both Chart-R1 Chen et al. (2025a) and BigCharts-R1 Masry et al. (2025) optimize MLLMs using Chain-of-Thought (CoT) Wei et al. (2022) reasoning data and reinforcement learning with verified rewards (RLVR) Shao et al. (2024).
@@ 2.2 Reasoning MLLMs
The success of large reasoning models (LRMs) like DeepSeek-R1 Guo et al. (2025) spurs significant research into enhancing LLM reasoning through reinforcement learning with rule-based rewards Shao et al. (2024). This paradigm extends to the vision-language domain, where numerous works aim to improve the long-chain reasoning capabilities of MLLMs Shen et al. (2025); Wang et al. (2025b). Following this trend, initial efforts such as Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025) adapt the Group Relative Policy Optimization (GRPO) algorithm with multimodal reasoning data to enable long-form reasoning in VLMs. Subsequent works, including MMEureka Meng et al. (2025b) and R1-Zero Liu et al. (2025), further advance this area by introducing improved reinforcement learning strategies for visual long-term reasoning. Recently, approaches like Point-RFT Ni et al. (2025) utilize task-specific thinking frameworks to enhance reasoning capabilities further.
@@ 2.3 MLLMs For Code
In the rapidly advancing research area of multimodal code generation, foundational work has focused on establishing benchmarks to assess the capabilities of MLLMs. These benchmarks evaluate performance on tasks such as generating execution code for solving visual problems Li et al. (2024); Wang et al. (2025a) and HTML code from webpage screenshots Si et al. (2024); Yun et al. (2024); Xiao et al. (2024). Among these emerging tasks, chart-to-code generation has attracted significant interest due to the complexity of its visual inputs. This task challenges an MLLM to generate plotting code that accurately reproduces the chart image. Recently, several benchmarks have been proposed to evaluate MLLMs in this context, assessing a range of capabilities. For instance, benchmarks such as ChartMimic Shi et al. (2024), Plot2Code Wu et al. (2024), and ChartX Xia et al. (2024) assess chart-to-code generation capabilities.
@@ 3.1 Training Corpus Construction
While several chart-to-code generation datasets have been proposed, they are generally constrained by a reliance on purely synthetic data and limited scale, resulting in charts that exhibit simplistic trends and lack diversity. To overcome these challenges, we construct our dataset by crawling real-world tables from arXiv papers and then leveraging LLMs to generate the corresponding plotting code. We ensure data diversity and quality through several strategies, such as using high-quality seed code and filtering out non-executable code.
Through this process, we construct the largest training corpus to date, comprising 3 million chart-code pairs, to investigate the performance limits of SFT. However, previous research Chen et al. (2025a); Huang et al. (2025) shows that separating the SFT and RL training corpora prevents overfitting to the SFT data format and enhances the exploration capacity of RL. To curate a high-quality dataset for RL, we begin with a candidate corpus of 130 thousand samples from the overall training corpus. These candidate data samples are then subjected to a two-stage filtering strategy. The first filtering stage focuses on the code content, filtering based on chart and data type to ensure code diversity and that the output code is suitable for data extraction. Specifically, the chart types are filtered according to plot functions. To handle complex charts containing multiple chart types, we employ a tree-structured parsing method to progressively identify each type present. The data values are filtered according to the data definition format, retaining only data structured as one-dimensional arrays and non-nested dictionaries to facilitate extraction during the RL stage. This step narrows the candidate corpus to 45 thousand samples. The second stage addresses visual quality, employing GPT-4.1-mini OpenAI (2025a) as a visual LLM judge to filter for high-fidelity chart images. This process yields the final RL training corpus of 33 thousand high-quality samples. The remaining samples are utilized for SFT.
@@ 3.3 Multimodal Structured Reinforcement Learning
Although SFT can significantly enhance the chart-to-code capabilities of MLLMs, our experiments reveal that this approach ultimately reaches a performance plateau, likely due to the limitation that SFT restricts the model’s capacity to restore detailed information. This limitation stems from the core drawback of the SFT objective that it treats every token in the target sequence with uniform importance. Therefore, SFT is ineffective in optimizing plotting code generation, as the plotting code comprises largely of boilerplate snippets such as plt.plot, with critical information like specific data values or styling parameters appearing infrequently. Consequently, employing RL with a reward function designed to prioritize the accuracy of such critical content presents a promising approach for enhancing code generation capacity. To this end, we propose a two-stage training strategy to stabilize the RL process, utilizing multi-granularity reward functions that combine both textual and visual feedback.
@@
Textual Reward While the RLVR method provides a powerful mechanism to guide models by measuring the correctness of detailed information, its practical application is challenged by the stylistic diversity of generated code, particularly in data definitions and function calls. This stylistic heterogeneity leads to difficulty in comprehensively extracting and parsing the critical information for reward computation. To mitigate this, we introduce a code normalization step that maps each generated output to a canonical representation before calculating the reward. This process ensures the textual reward function is invariant to syntactic variations, providing a more suitable method for RLVR. Specifically, the textual reward is multifaceted, consisting of a binary score for execution success and a more granular rule-based accuracy score. This accuracy metric is a weighted average derived from evaluating key aspects of the generated code. We assess data values using soft value matching, chart types with hard string matching, layout with hard value matching and elements like titles and labels via edit distance. The soft value matching allows a relative error tolerance of $\pm$5%. The final textual reward is the addition of execution and a rule-based accuracy score.
@@
Visual Reward However, our analysis reveals that purely textual rewards generally focus on fine-grained details, while the overall structure of the rendered chart image is not considered. The chart-to-code task requires generating code that replicates the overall style of charts. To this end, we propose visual reward feedback. Specifically, the generated code is executed to render a chart image, and we then utilize MLLMs to score the visual similarity between the generated chart and the input. This score is then normalized to serve as the visual reward. For the visual reward, we employ Qwen2.5-VL-72B Bai et al. (2025) as the evaluation model, which compares the chart image rendered by the generated code against the ground-truth chart across six key aspects: chart type, layout, text content, data, style, and clarity. The resulting scores are then normalized to form the final reward. Crucially, code that fails to render an image receives a reward of 0. The framework of our proposed MSRL is denoted in Figure 2.
@@
Two-stage RL To ensure training stability, our two-stage RL strategy first optimizes the model using only a textual reward. In the second stage, we introduce a hybrid reward, which combines the textual reward with a visual one, to fine-tune the model for visual fidelity. To properly balance the different training objectives, the total reward $R$ for a completed trajectory is calculated as a weighted sum of the following three components:
@@
where $w_{t}$, $w_{v}$, and $w_{e}$ are hyperparameters that balance the contribution of each component.
GRPO We adapt the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm for RL. For each input $(x,y)$, the policy $\pi_{\theta}$ samples a group of $G$ candidate responses $\{o_{i}\}_{i=1}^{G}$.
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{\begin{subarray}{c}(x,y)\sim\mathcal{D}_{\text{CoT}},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid x)\end{subarray}}$ $\biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\biggl{(}\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)}A_{i},$ (2)
$\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)},1-\varepsilon,1+\varepsilon\right)A_{i}\biggr{)}]$
where $\varepsilon$ is a hyperparameter, and $\pi_{\theta}$, $\pi_{\theta_{\text{old}}}$ are the optimized model and the old policy model. The group-normalized advantage for the $i$-th response is:
$A_{i}=\frac{r_{i}-\operatorname{mean}\left(\left\{r_{1},r_{2},\cdots,r_{G}\right\}\right)}{\operatorname{std}\left(\left\{r_{1},r_{2},\cdots,r_{G}\right\}\right)}$ (3)

### v2

@@ Abstract
While reinforcement learning (RL) has proven highly effective for general reasoning in vision-language models, its application to tasks requiring deep understanding of information-rich images and structured output generation remains underexplored. Chart-to-code generation exemplifies this challenge, demanding complex reasoning over visual charts to produce structured code. Supervised fine-tuning (SFT) alone is often insufficient, highlighting the need for effective RL strategies tailored to structured outputs. In this paper, we systematically investigate the performance plateau of SFT through large-scale experiments and propose Multimodal Structured Reinforcement Learning (MSRL) for chart-to-code generation. We construct the largest training corpus to date, with 3 million chart-code pairs curated from real-world tables in arXiv papers, addressing the limitations of previous synthetic datasets. Despite achieving state-of-the-art performance, our experiments show that simply increasing SFT data eventually leads to diminishing improvements. To break this plateau, MSRL employs a multi-granularity reward system that integrates both textual and visual feedback. At the textual level, rule-based rewards validate fine-grained code details, while at the visual level, a model-based reward assesses the structural similarity between rendered code and ground-truth charts. We implement a two-stage curriculum training strategy, first optimizing the model with textual rewards and then incorporating visual signals for further enhancement. Experimental results demonstrate that MSRL substantially breaks the SFT plateau, improving high-level metrics by 6.2% and 9.9% on ChartMimic and ReachQA benchmarks, respectively. Notably, our method outperforms all existing approaches in the chart domain and achieves competitive results with advanced closed-source models.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive reasoning capabilities on complex textual problems, including code generation and mathematical problem-solving OpenAI (2025b); Guo et al. (2025). Pivotal to this success is the implementation of reinforcement learning (RL) to optimize the capacities of LLM to generate step-by-step reasoning Wei et al. (2022) and verify final answers Shao et al. (2024). While multimodal large language models (MLLMs) have achieved impressive performance on established visual reasoning tasks like visual question answering Zhang et al. (2024a); Huang et al. (2025), their core reasoning capabilities remain limited, especially when tasks require processing information-dense images like charts Xu et al. (2024); Chen et al. (2025a), and generating structured outputs like code Xia et al. (2024); Zhao et al. (2025b).
Among these challenges, chart-to-code generation Yang et al. (2024a); Zhang et al. (2024c) stands out as a task of both high complexity and significant practical value. Charts are widely used to communicate scientific findings, business insights, and statistical data, yet their information is often locked in visual formats that are difficult for automated systems to interpret or repurpose. Enabling MLLMs to accurately convert charts into executable code not only advances their visual understanding and structured reasoning capabilities, but also opens up impactful applications, such as assisting researchers in automating the reproduction of scientific visualizations. Although previous works have attempted to solve this task by using methods like Supervised Fine-Tuning (SFT) Zhang et al. (2024b); Zhao et al. (2025b) or Direct Preference Optimization (DPO) Rafailov et al. (2023); Zhang et al. (2025) on synthetic data, their performance remains limited due to the simplistic patterns of the corpus and ineffective generalization strategies. Therefore, a promising direction is to leverage more realistic data and robust strategies to enhance model performance.
@@
In this work, we first systematically investigate the SFT performance plateau through experiments at various data scales. To enable this analysis, we construct the largest training corpus to date for chart-to-code generation, covering 24 distinct chart types and comprising 3 million chart-code pairs. Our two-stage data pipeline curates high-quality tables from arXiv papers and then utilizes LLMs to generate diverse and complex plot codes, ensuring realistic distributions and avoiding monotonous trends. Unlike prior works Ni et al. (2025); Chen et al. (2025b) that apply RL directly after SFT, we establish saturated SFT baselines to isolate the true effectiveness of our proposed multimodal structured reinforcement learning (MSRL) strategy. Through experiments with varying SFT data scales, we demonstrate that simply increasing the volume of SFT data leads to a performance plateau that cannot be overcome by data quantity alone. To break through this plateau, we propose the MSRL training strategy, which incorporates multi-granularity reward functions from both textual and visual aspects. For textual rewards, we standardize code formats in the training corpus and design a comprehensive reward function that verifies code correctness from five key perspectives. Recognizing that text-based rewards alone may overlook overall chart structure, we introduce a visual reward mechanism: generated code is rendered into images and an MLLM evaluates their similarity to the original input chart. By jointly leveraging these multi-granularity rewards, MSRL optimizes both global chart context and fine-grained code details. We evaluate models on various chart-to-code benchmarks Yang et al. (2024a); He et al. (2024). The results demonstrate that it establishes a new state-of-the-art, as it outperforms all previous open-source models and rivals advanced proprietary MLLMs. In summary, the main contributions of this work are as follows:
@@ 2.1 Chart MLLMs
Chart understanding is a crucial research area encompassing both low-level tasks, such as data extraction Liu et al. (2023), and high-level tasks like question answering Masry et al. (2022) and summarization Kantharaj et al. (2022). Recent work has focused on training MLLMs on extensive, chart-specific datasets to enhance their understanding capabilities Xia et al. (2024); Zhang et al. (2024b); Xia et al. (2023), leading to superior performance across various chart-related tasks. For example, ChartMoE Xu et al. (2024) proposes a mixture-of-experts (MoE) structure for multi-task pretraining. More recently, leveraging reinforcement learning Chen et al. (2025a); Masry et al. (2025); Rodriguez et al. (2025b) to enhance reasoning capacities has garnered significant interest. Both Chart-R1 and BigCharts-R1 optimize MLLMs using Chain-of-Thought (CoT) Wei et al. (2022) reasoning data and reinforcement learning with verified rewards (RLVR) Shao et al. (2024).
@@ 2.2 Reasoning MLLMs
The success of large reasoning models (LRMs) like DeepSeek-R1 Guo et al. (2025) spurs significant research into enhancing LLM reasoning through reinforcement learning with rule-based rewards Shao et al. (2024). This paradigm extends to the vision-language domain, where numerous works aim to improve the long-chain reasoning capabilities of MLLMs Shen et al. (2025); Wang et al. (2025b). Following this trend, initial efforts such as Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025b) adapt the Group Relative Policy Optimization (GRPO) algorithm with multimodal reasoning data to enable long-form reasoning in VLMs. Subsequent works, including MMEureka Meng et al. (2025) and R1-Zero Liu et al. (2025), further advance this area by introducing improved reinforcement learning strategies for visual long-term reasoning. Recently, approaches like Point-RFT Ni et al. (2025) utilize task-specific thinking frameworks to enhance reasoning capabilities further.
@@ 2.3 MLLMs For Code
In the rapidly advancing research area of multimodal code generation, foundational work has focused on establishing benchmarks to assess the capabilities of MLLMs. These benchmarks evaluate performance on tasks such as generating execution code for solving visual problems Li et al. (2024); Wang et al. (2025a); Rodriguez et al. (2025a; 2024); Yang et al. (2024b) and HTML code from webpage screenshots Si et al. (2024); Yun et al. (2024); Xiao et al. (2024); Awal et al. (2025). Among these emerging tasks, chart-to-code generation has attracted significant interest due to the complexity of its visual inputs. This task challenges an MLLM to generate plotting code that accurately reproduces the chart image. Recently, several benchmarks have been proposed to evaluate MLLMs in this context, assessing a range of capabilities. For instance, benchmarks such as ChartMimic Yang et al. (2024a), Plot2Code Wu et al. (2025), and ChartX Xia et al. (2024) assess chart-to-code generation capabilities, while others like ChartEdit Zhao et al. (2025a) and $\text{ChartM}^{\text{3}}$ Yang et al. (2025a) evaluate the editing functionalities.
@@ 3.1 Training Corpus Construction
While several chart-to-code generation datasets have been proposed, they are generally constrained by a reliance on purely synthetic data and limited scale, resulting in charts that exhibit simplistic trends and lack diversity. To overcome these challenges, we construct our dataset by crawling real-world tables from arXiv papers and then leveraging LLMs to generate the corresponding plotting code. To ensure separation from the ChartMimic benchmark, we exclusively use papers published in 2023 and earlier as our corpus. We ensure data diversity and quality through several strategies, such as using high-quality seed code and filtering out non-executable code.
Through this process, we construct the largest training corpus to date, comprising 3 million chart-code pairs, to investigate the performance limits of SFT. However, previous research Chen et al. (2025a); Huang et al. (2025) shows that separating the SFT and RL training corpora prevents overfitting to the SFT data format and enhances the exploration capacity of RL. To curate a high-quality dataset for RL, we begin with a candidate corpus of 130 thousand samples from the overall training corpus. These candidate data samples are then subjected to a two-stage filtering strategy. The first filtering stage focuses on the code content, filtering based on chart and data type to ensure code diversity and that the output code is suitable for data extraction. Specifically, the chart types are filtered according to plot functions. To handle complex charts containing multiple chart types, we employ a tree-structured parsing method to identify each type present progressively. The data values are filtered according to the data definition format, retaining only data structured as one-dimensional arrays and non-nested dictionaries to facilitate extraction during the RL stage. This step narrows the candidate corpus to 45 thousand samples. The second stage addresses visual quality, employing GPT-4.1-mini OpenAI (2025a) as a visual LLM judge to filter for high-fidelity chart images. To mitigate potential bias from the MLLM judger, we manually inspect 100 samples that are assigned a score of 0 by the model, and observe over 90% consistency with human judgment. This process yields the final RL training corpus of 33 thousand high-quality samples. The remaining samples are utilized for SFT. The comparisons of our dataset with relevant chart-to-code datasets are listed in Table 1. Our dataset features a completely realistic data source, more challenging multi-chart content, and richer visual diversity enabled by more API types.
Table 1: Comparison of existing chart-to-code generation datasets. ✔
indicates that ChartMoE uses both data from real CSV tables and data generated by LLMs. ”API types” refers to the number of different Matplotlib API types covered by the dataset.
Dataset Realistic data source Multi-chart Chart types Data samples API types
ChartLlama 10 11k 83
ChartMoE
<<20 800k -
Chart2Code 15 3k 168
ChartCoder 27 115k 187
Ours 24 3M 1,555
@@ 3.3 Multimodal Structured Reinforcement Learning
Although SFT can significantly enhance the chart-to-code capabilities of MLLMs, our experiments reveal that this approach ultimately reaches a performance plateau, likely due to the limitation that SFT restricts the model’s capacity to restore detailed information. This limitation stems from the core drawback of the SFT objective that it treats every token in the target sequence with uniform importance. Therefore, SFT is ineffective in optimizing plotting code generation, as the plotting code comprises largely of boilerplate snippets such as plt.plot, with critical information like specific data values or styling parameters appearing infrequently. Consequently, employing RL with a reward function designed to prioritize the accuracy of such critical content presents a promising approach for enhancing code generation capacity. To this end, we propose a two-stage RL training strategy utilizing multi-granularity reward functions that combine both textual and visual feedback.
@@
Textual Reward While the RLVR method provides a powerful mechanism to guide models by measuring the correctness of detailed information, its practical application is challenged by the stylistic diversity of generated code, particularly in data definitions and function calls. This stylistic heterogeneity leads to difficulty in comprehensively extracting and parsing the critical information for reward computation. To mitigate this, we introduce a code normalization step that maps each generated output to a canonical representation before calculating the reward. This process ensures the textual reward function is invariant to syntactic variations, providing a more suitable method for RLVR. Specifically, the textual reward is a granular rule-based accuracy score, which is a weighted average derived from evaluating key aspects of the generated code. We assess data values using soft value matching, chart types with hard string matching, layout with hard value matching and elements like titles and labels via edit distance. The soft value matching allows a relative error tolerance of $\pm$5%. Separately, we compute an execution reward, which is a binary score indicating whether the generated code can be executed successfully. These two rewards ensure that both the fidelity of critical information and the executability of the code are considered during RL optimization.
@@
Visual Reward However, our analysis reveals that purely textual rewards generally focus on fine-grained details, while the overall structure of the rendered chart image is not considered. The chart-to-code task requires generating code that replicates the overall style of charts. To this end, we propose visual reward feedback Gu et al. (2025). Specifically, the generated code is executed to render a chart image, and we then utilize MLLMs to score the visual similarity between the generated chart and the input. This score is then normalized to serve as the visual reward. For the visual reward, we employ Qwen2.5-VL-72B Bai et al. (2025) as the evaluation model, which compares the chart image rendered by the generated code against the ground-truth chart across six key aspects: chart type, layout, text content, data, style, and clarity. The resulting scores are then normalized to form the final reward. Crucially, code that fails to render an image receives a reward of 0. The framework of our proposed MSRL is denoted in Figure 2.
@@
Two-stage RL Obtaining visual rewards through model evaluation requires substantial computational time. To ensure training efficiency, our two-stage RL strategy first optimizes the model using only a textual reward. In the second stage, we introduce a hybrid reward, which combines the textual reward with a visual one, to fine-tune the model for visual fidelity using a reduced number of samples. This approach allows us to balance computational cost and model performance, achieving strong visual quality without excessive resource consumption. To properly balance the different training objectives, the total reward $R$ for a completed trajectory is calculated as a weighted sum of the following three components:
@@
where $w_{t}$, $w_{v}$, and $w_{e}$ are hyperparameters that balance the contribution of each component. We adopt the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm for RL. GRPO optimizes policy by leveraging group-wise relative advantages among sampled responses, which is well-suited for our customized reward designs of charts.